OpenStudy (anonymous):

The amounts of paper waste generated in a region during two years were 41,300,000 tons and 50,500,000 tons. What was the total paper waste generated over these two years in the region, expressed in scientific notation?

OpenStudy (gschibby):

That would be \[Total amount = 41,3*10^{6}+50,5*10^{6}=?\]

OpenStudy (anonymous):

Add these two values and represent that value in 10^x

OpenStudy (anonymous):

Simplify (36)18.

OpenStudy (gschibby):

(36)18=36*18=?